OpenStudy (anonymous):

Simplify. sqrt108 - 3sqrt8 - sqrt27

OpenStudy (zepp):

Umm, do you know that rule about if you have something like \(\sqrt{9}\) you can split it into \(\sqrt{3*3}\) ?

OpenStudy (zepp):

Similary, you can split \(\sqrt{162}\) into \(\sqrt{9*9*2}\) hence \(9\sqrt{2}\), can you do that for everything you have there?

OpenStudy (precal):